# biblio.bib

### v1

@@ Abstract
Mechanical forces are crucial regulators of diverse biological processes, yet their in vivo measurement remains challenging due to limitations of current techniques, that can be destructive or require complex dedicated setups. This study introduces a novel class of biocompatible, self-functionalizing stress sensors based on inverted emulsions, that can be used to measure mechanical stresses inside tissues but can also locally perturb the biological environment through specific binder presentation or drug delivery. We engineered an optimal design for these inverted emulsions, focusing on finding the perfect balance between the two contradictory constraints: achieving low surface tension for deformability while maintaining emulsion instability for efficient self-functionalization and drug release. Proof-of-concept experiments in both agarose gels and complex biological systems, including brain organoids and zebrafish embryos, confirm the droplets ability to deform in response to mechanical stress applied on the tissue, to self-functionalize and to release encapsulated molecules locally. These versatile sensors offer a method for non-invasive stress measurements and targeted chemical delivery within living biological tissues, overcoming current technical barriers in biophysical measurements.
@@
emulsions, stress measurements, functionalization, drug release1 Introduction
@@
It is well-established that mechanical forces bear a central role during many biological processes such as morphogenesis or cancer progression. At the tissue-scale, these forces can be exerted intrinsically, through the motile and contractile activity of the cells within a tissue, or can be sustained from externally applied forces, for instance through the activity of neighboring tissues or through mechanical coupling via the extracellular matrix.
@@
A lot of effort has therefore been dedicated in the past decades to map out these forces and the associated mechanical properties of tissues undergoing morphogenetic processes. For instance, various bulk rheometry or tensiometry techniques have been adapted to the use of biological material [Mary2022, Mazuel2015, Stirbat2013, Mgharbel2009, Foty1994]. Micromanipulation assays and microfluidics were also developed to gain insight into the mechanical properties of tissues at different scales [Guevorkian2010, Tlili2022]. However, these techniques were mostly applied to cellular aggregates or surface tissues [vonDassow2010, Petridou2021] in living organisms because of the constraints of set-up geometry .
@@
Other tools can be used to infer local mechanical properties and stresses in vivo, and at different scales. Among these tools, laser ablation is a powerful technique allowing to map out tension within developing tissues [Rauzi2008, Breau2017, Porazinski2015]. However, its destructive nature imposes to renew the experiment on multiple organisms in order to map tension across space and time. Moreover, an absolute measurement of tension requires to know the material properties of the destructed area. Alternatively, sensors in the form of gel beads or droplets have been developed in the past few years, which allows to measure mechanical properties in vivo with a non-destructive approach (see [Ding2022] for a review). Compressible gel beads inserted within aggregates or living tissues thus allow to measure not only compressive [Dolega2017, Lee2019, Vian2023], but also tensile stresses when the sensors are coupled to their biological environment [Mohagheghian2018, Traber2019, Gutierrez2021]. On the other hand, incompressible but deformable sensors allow for the measurement of anisotropic stresses generated inside tissues [Souchaud2022, Campas2014, Mongera2018].
Here again, specific adhesion between the droplets and their environment can be implemented [Campas2014], which is necessary in order to unravel the role of adhesive forces during biological processes. However, such adhesive droplets have not been used extensively in further studies. This might be due to the fact that droplets were preferentially formed in situ through microinjection rather than injected directly as a dispersed emulsion, which is necessary in the case of functionalized adhesive droplets.
@@
In this paper, we introduce a new type of stress measurement sensors which can gain secondary functions once injected inside the tissue, such as local drug delivery or surface exposure of proteins. This so-called self-functionalization in situ is due by the fact that the droplet sensors are made of an inverted emulsion of water droplets in oil. The inverted emulsion is designed to be unstable enough so that the inner droplets progressively fuse at the oil-water outer surface, thus releasing locally any drug that was initially contained inside the droplets or grafted onto their surface. We report how to fine tune the balance between surface tension - controlling self-functionalization - and emulsion stability - controlling drug delivery - so that both functions of the sensors can take place inside biological materials such as soft developing tissues. Moreover, this technique allows one to inject single droplets inside the tissue and to precisely localize them. All of these characteristics confer to these sensors a great versatility of functions and applicability in both inert and biological materials.
@@ 2.1 Design of the inverted emulsion sensor
A variety of techniques have been developed to produce biomimetic functional oil droplets [Pontani2012]. These systems were mostly used to mimic cell-cell adhesion [Pontani2012, Nagendra2023], allowing them to interact specifically with cells or biological structures to unravel biological processes such as phagocytosis [BenMBarek2015] or intracellular transport [Bottier2009]. In these cases, the droplets are initially designed in the form of an oil in water emulsion that is afterwards injected or delivered to a specific location. The emulsions can be made through a variety of techniques, such as controlled shear in a Couette apparatus [Pontani2012, BenMBarek2015, Pinon2018], extrusion through porous membranes [Pontani2013, Zhang2017] or microfluidics [Nagendra2023]. In their dispersed form they can be used to measure forces associated with single cell processes, in which they can simply be put in the presence of cells inside dedicated set-ups. In vivo, the use of such emulsions implies that a mixture of oil droplets and water has to be injected directly inside the tissues. Although this approach has been applied in explants and cellular aggregates [Campas2014, Souchaud2022],injecting such complex materials inside developing tissues remains challenging. The water phase contained in the emulsion may swell the tissue locally, leading to a large stress and potential exclusion of the droplets from the tissue. Moreover, the dispersed and inherently heterogeneous nature of the emulsion make it more difficult to inject single droplets at a precise locations. Therefore, most of the in vivo studies using oil based sensors actually turned towards the use of single droplets that are microinjected from a bulk oil phase [Mongera2018].
In order to combine formulation versatility and ease of the injection process, we thus explored the use of inverted emulsions as a base for the microinjection of isolated droplets. Upon destabilization of the inverted emulsion, the inner water droplets should fuse with the outer oil water interface, thus releasing any chemical that is encapsulated inside them (Fig.1A). Moreover, if the inner surface of the droplets is functionalized with proteins or active molecules - that we refer to as binders in the following - the outer surface of the droplets can become decorated with the same binders upon fusion (Fig.1B), thus providing self-functionalizing properties to the injected droplet.
In order for this schematic to work, the inverted emulsion has to be unstable. This means that there should be just enough surfactant in the oil to ensure potential functionalization, but not too much of it, which would tend to stabilize the inner water droplets and impair both drug release and relocalization of the functionalization to the outer surface. In the meantime, the droplet needs to be deformable for stress measurements. To this end, there should be enough surfactant in the oil to lower the surface tension of the oil/water interface.
In the following, we therefore describe the formation and composition of the inverted emulsion, as well as the systematic measurements of surface tension and destabilization efficiency. These compared measurements allowed us to identify a sweet spot for formulation in which both properties can be fulfilled.
@@ 2.2 Formation of the inverted emulsion
The inverted emulsion was first made of water droplets dispersed in a biocompatible soybean oil containing various concentrations of phospholipids and Tween 20, both of which act as biocompatible surfactants. The chosen lipid, called 16:0 Biotinyl PE, carries a biotin on its hydrophilic head. This allows us to visualize the surface of the inner droplets by encapsulating fluorescent streptavidin in the water phase, which gets spontaneously grafted onto the biotinylated lipids heads (Fig.1C). Moreover, the structure of this biotinylated lipid is compatible with an inverted emulsion topology as its hydrophilic head is smaller compared to the classical biotinylated lipids used for direct oil in water emulsions, such as DSPE-PEG(2000) Biotin lipids whose hydrophilic head is about five times larger than that of 16:0 Biotinyl PE.
@@
The dispersion of the aqueous phase can be performed with different methods depending on the material to be encapsulated inside the inverted emulsion. When using expensive and rare biological material (such as antibodies or purified proteins), the emulsification is adapted to the use of small volumes. In that case, of aqueous phase is gently emulsified in the oil phase through a porous Shirasu Porous Glass (SPG) membrane (see Fig. 2 and Materials and Methods). This emulsification method only requires very little material (the SPG connector itself and two syringes) and can be performed manually within seconds. This is therefore perfectly adapted to the use of fragile biological material and can be performed straightforwardly in any laboratory environment.
With a membrane of pore size , this method typically yields an emulsion with a mean diameter of a standard deviation of using our formulation with lipid and Tween 20 (see Fig 3 and Methods).
@@
From the droplet size distribution, we can estimate the quantity of surfactant required to cover the typical surface area of oil-water interface in the emulsion. Indeed, the size distribution leads to a mean volume of for the water droplets, while the water volume fraction , with and yields (see Methods). Considering the resulting total surface area in the emulsion and assuming a typical surface coverage of lipids , for the chosen lipids (16:0 Biotinyl PE, ) this leads to a lipid mass concentration in the oil of to cover the inner water droplets. Because the lipids do not all partition at the oil/water interface, we therefore work at lipid concentrations .
@@
The binders concentration also needs to be adapted to the amount of available lipids on the surface. In order to avoid the presence of unattached binders in the droplets, that could be freely released in the surrounding medium, their concentration must be lower than that of the surface lipids. Moreover, as mentioned above, not all lipids will partition at the oil/water interface, especially in the presence of co-surfactants such as Tween. Therefore we work at a ratio of about 1 streptavidin for 2000 lipids. Considering the size of the water droplets in the inverted emulsion, this target value leads to a streptavidin concentration in the droplets of .
@@
If the material contained in the water droplets is more robust and abundant, the emulsification method can be simply performed by pushing the oil and water mix through a classical PVDF (Polyvinylidene fluoride) filter with a pore size (see Materials and Methods). The dead volume with these filters is quite important, which is why it is not adapted to the handling of precious material such as purified proteins, but can be used for classical fluorophores or drugs. For the same formulation ( lipid and Tween 20), we get an emulsion that is a bit coarser with the filter compared to the SPG membrane, with a mean diameter of and a standard deviation of . With an alternative formulation containing only lipid, we get an even coarser emulsion, with a mean diameter of a standard deviation of , which is expected since the surface tension of this formulation is higher (see Fig.5E).
@@ 2.3 Tuning of the interfacial tension
In order to use droplets as stress sensors, their interfacial tension must low enough and is therefore a critical parameter of our design. Indeed, when injected in a soft biological tissue or a gel, an oil droplet can initially adopt a spherical shape minimizing its area of contact with the surrounding material. If the material is deformed, the droplet will in turn exhibit a measurable deformation given that its interfacial tension is low enough, as detailed below.
@@
As an illustration, we consider a droplet injected in a surrounding medium that is in steady state under uniaxial compression . We thus consider the surrounding medium as an elastic material, with an associated Young’s modulus . In this representation, illustrated in Fig. 4, the droplet is under an elastic stress that depends on the surrounding material Young’s modulus [sunMechanicsFibrillarCollagen2021]. The capillary stress , which linearly depends on the interfacial tension of the oil/water interface, is resisting droplet deformation. In order to get an effective stress sensor, capillary stress and elastic stress should have the same order of magnitude. If the elastic stress is much bigger than capillary stress, the droplet will be under a large deformation regime, up to break up. If the capillary stress is much bigger than the elastic stress, the droplet will not be deformed, thus behaving as a solid inclusion in a soft material. In intermediate regimes, deformations of the included droplet can inform on local stresses, given that the Young modulus and surface tension are known [Tapie2023]. In soft developing tissues or biopolymers such as extracellular matrices, the material Young’s modulus typically falls in the range [bohringerDynamicTractionForce2024]. The droplet size is constrained by the one of the considered biological tissue and the typical length scale of the process being probed. Typically, droplets of radius are usually small enough to probe cellular scale processes in tissues and large enough to be imaged with a sufficient resolution. We therefore use this minimal size to deduce the required maximal interfacial tension for observable droplet deformations, which yields:
@@
To assess if such values are accessible in our system, we first tested the effect of lipids alone on interfacial tension. To this end, we measured the interfacial tension through a rising drop method (see Fig. 5). Importantly, we also measured the interfacial tension in between an inverted emulsion and a continuous aqueous phase, which matches more accurately our system’s configuration.
@@
As the concentration of lipids is increased from to in the oil, the interfacial tension decreases. Notably, the same concentrations of lipids lead to similar interfacial tensions with a rising drop made of an inverted emulsion (see Fig. 5), thus validating our system for stress measurements in vivo. However, in both cases, the tension does not go beyond , even at lipids. Yet, at such lipid concentrations we expect the inner emulsion to be quite stable [Pontani2012] and to hinder the targeted self-functionalizing features of the sensor.
@@
To overcome this limitation, the use of Tween 20 as a co-surfactant is critical. Indeed, it is a synthetic surfactant that is largely used to stabilize direct oil in water emulsions. Therefore, we expect that the addition of Tween 20 should preferentially lower the oil/water interface at the outer surface of the sensor (positive curvature), while having a more marginal effect on the stability of the inner water droplets whose negative curvature does not favor Tween insertion. To test the effect of Tween 20 on surface tension, we measured the oil/water interfacial tension for oils containing a fixed concentration of lipids at and Tween 20 concentrations up to (see Fig. 5).
@@
Strikingly, an addition of Tween 20 already lowers the interfacial tension down to , for both the oil and inverted emulsion systems. Increasing the Tween 20 concentration up to only marginally lowers the tension down to and could again hinder the destabilization process. However, in both cases, this lowered tension is in principle low enough to measure stresses inside soft tissues as discussed above. In the following, we explore the efficiency of the inverted emulsion destabilization and self-functionalization of injected droplets, as a function of surfactant concentrations and composition.

### v3

@@ Abstract
Mechanical contributions are crucial regulators of diverse biological processes, yet their in vivo measurement remains challenging due to limitations of current techniques, that can be destructive or require complex dedicated setups. This study introduces a novel method to synthesize biocompatible, self-functionalizing stress sensors based on inverted emulsions, that can be used to probe stresses inside tissues but can also locally perturb the biological environment through specific binder presentation or drug delivery. We engineered an optimal design for these inverted emulsions, focusing on finding the balance between the two contradictory constraints: achieving low surface tension for deformability while maintaining emulsion instability for efficient self-functionalization and drug release. Proof-of-concept experiments in both agarose gels and complex biological systems, including brain organoids and zebrafish embryos, confirm the droplets ability to deform in response to mechanical stress applied within the tissue, to self-functionalize and to release encapsulated molecules locally. These versatile sensors offer a method for non-invasive stress measurements and targeted chemical delivery within living biological tissues, giving the potential to overcome current technical barriers in biophysical studies.
@@
oil droplets, stress sensors, self-functionalization1 Introduction
@@
It is well-established that mechanical forces bear a central role during many biological processes such as morphogenesis [1, 2, 3] or cancer progression [4]. At the tissue-scale, these forces can be exerted intrinsically, through the motile and contractile activity of the cells within a tissue, or can be sustained from externally applied forces, for instance through the activity of neighboring tissues or through mechanical coupling via the extracellular matrix [5].
@@
A lot of effort has therefore been dedicated in the past decades to map out these forces and the associated mechanical properties of tissues undergoing morphogenetic processes (see [6, 7] for reviews). In vitro, various bulk rheometry or tensiometry techniques have been adapted to the use of biological material [8, 9, 10, 11, 12]. Micromanipulation assays and microfluidics were also developed to gain insight into the mechanical properties of tissues at different scales [13, 14]. However, these techniques were mostly applied to cellular aggregates or surface tissues [15, 16] in living organisms because of the constraints of set-up geometry.
@@
Other tools can be used to infer local mechanical properties and stresses in vivo, and at different scales. Among these tools, laser ablation is a powerful technique allowing to map out tension within developing tissues [17, 18, 19, 20]. However, its destructive nature imposes to renew the experiment on multiple organisms in order to map tension across space and time. Moreover, an absolute measurement of tension requires to know the material properties of the destructed area.
Alternatively, sensors in the form of deformable inclusions - encompassing solid beads and liquid droplets - have been developed in the past few years, and have been used to measure mechanical stresses in vivo with a non-destructive approach (see [21] for a review). On the one hand, solid beads composed of compressible gels offer several advantages. Their mechanical compressibility can be precisely calibrated under controlled conditions (for instance, by applying osmotic pressure), enabling the measurement of isotropic stresses within biological tissues such as cell aggregates or developing embryos [22, 23, 24]. In addition, these beads can be functionalized to interact specifically with their surrounding environment, making them a versatile tool for probing not only compressive but also tensile stresses [25, 26, 27].
On the other hand, incompressible but deformable sensors allow for the measurement of anisotropic stresses generated inside the tissues [28, 29, 30]. In particular, oil droplets can be functionalized with binders [31, 32, 33, 34] in order to induce specific adhesion between the droplets and their environment. In this case, the oil droplets are prepared and functionalized prior to their injection inside the tissues, in the form of an emulsion. This approach has been applied in explants and cellular aggregates [29, 35], but remains limited for in vivo studies. Indeed, for both solid and liquid sensors, the use of such dispersed solutions (emulsions for liquid sensors, dispersions of beads for solid ones) implies that a mixture of sensors and water has to be injected directly inside biological tissues, which remains challenging for multiple reasons. First, the water phase contained in the solution may swell the tissue locally, leading to a large stress and potential tissue damage and exclusion of the sensors from the tissue. Second, the dispersed and inherently heterogeneous nature of the dispersions make it very difficult to inject a limited number of sensors at a precise location. Therefore, in vivo studies actually turned towards the microinjection of single oil droplets from a pure oil phase [30], thus loosing the surface functionalization properties of classical biomimetic emulsions but gaining in precision and biocompatibility regarding the injection itself.
@@
In this paper, we introduce a new method to synthesize stress measurement sensors based on oil droplets which can gain secondary functions once injected inside the tissue, such as local drug delivery or surface exposure of proteins. This so-called self-functionalization in situ is due to the fact that the droplet sensors are made of an inverted emulsion of water droplets in oil. The inverted emulsion is designed to be unstable enough so that the inner water droplets progressively fuse at the oil-water outer surface, thus releasing locally any drug that was initially contained in the inner water droplets or exposing any protein grafted on their surface. In order for the inverted emulsion to be unstable, there should therefore be a limited amount of surfactants in the formulation. On the other hand, for oil droplets to act as stress sensors, i.e. be deformed in biological tissues, their size has to be at least comparable to the elasto-capillary length which is given by , with the interfacial tension of the oil droplet and the young modulus of the surrounding tissue. If the condition on the elasto-capillary length is fulfilled, the included oil droplets can be deformed by their environment and these deformations can in turn provide a measurement of local stresses. Therefore, the smaller the surface tension, the smaller the constraint on the oil droplet size, making it a critical parameter for the formulation of stress sensors.
We here present a proof of concept study for the use of such oil droplets as functional stress sensors. We describe how to fine tune the balance between self-functionalizing properties - requiring inverted emulsion instability, i.e. a high enough surface tension - and stress sensor properties - requiring low enough surface tension - so that both functions of the sensors can take place inside biological materials such as soft developing tissues. We then use a mechanical excitation set-up to validate a mechanical model of such liquid insertions inside model materials. Finally, since this technique allows one to inject single oil droplets of a chosen size inside the tissue and to precisely localize them, we present a few examples of their incorporation in different materials, from agar gels to zebrafish embryos.
@@ 2.1 Design of the inverted emulsion sensor
In order to combine formulation versatility and ease of the injection process, we explored the idea of using inverted emulsions as a base for the microinjection of isolated oil drops inside tissues. Upon destabilization of the inverted emulsion, the inner water droplets should fuse with the outer oil water interface, thus releasing any chemical that is encapsulated inside them (Fig.1A). Moreover, if the inner surface of the water droplets is functionalized with proteins or active molecules - that we refer to as binders in the following - the outer surface of the drop can become decorated with the same binders upon fusion (Fig.1B), thus providing it with self-functionalizing properties.
In order for this schematic to work, the inverted emulsion has to be unstable. On the one hand there should therefore not be too much surfactant, which would tend to stabilize the inner water droplets. On the other hand, for the oil droplet to be deformable in biological tissues, there should be enough surfactant in the oil to lower its surface tension. In the following, we describe the formation and composition of the inverted emulsion, as well as the systematic measurements of surface tension and destabilization efficiency. These compared measurements allowed us to identify a sweet spot for formulation in which both properties can be fulfilled.
@@ 2.2 Formation of the inverted emulsion
The inverted emulsion is made of water droplets dispersed in a biocompatible soybean oil containing various concentrations of phospholipids and Tween 20, both of which act as biocompatible surfactants. The chosen lipid, called 16:0 Biotinyl PE, carries a biotin on its hydrophilic head. This allows us to visualize the surface of the inner water droplets by encapsulating fluorescent streptavidin in the water phase, which gets spontaneously grafted onto the biotinylated lipids heads (Fig.1C). This lipid was chosen because of its Hydrophilic-Lipophilic Balance index (HLB). This value can be estimated from the size of hydrophilic and hydrophobic domains of a surfactant and provides insight into the structure of a surfactant and its ability to stabilize rather direct (8HLB20) or inverse emulsions (HLB6) [36]. For 16:0 Biotinyl PE, the HLB index yields 5-6, which is low enough to stabilize water in oil emulsions but still quite hydrophilic. As a point of comparison, a DSPE-PEG(2000) Biotin lipids, that are commonly used to functionalize oil droplets [31], have an HLB index of 12-15, which would not be appropriate to populate the surface of the inner water droplets.
2.2.1 Emulsification through SPG membranes
@@
The dispersion of the aqueous phase can be performed with different methods. When using expensive and rare biological material (such as antibodies or purified proteins), the emulsification is adapted to the use of small volumes. In that case, of aqueous phase is gently emulsified in the oil through a porous Shirasu Porous Glass (SPG) membrane (see Fig. 2 and Methods). This emulsification method only requires very little material (the SPG connector itself and two syringes) and can be performed manually within seconds. It is therefore perfectly adapted to the use of fragile biological material and can be implemented straightforwardly in any laboratory environment. With a membrane of pore size , this method typically yields an emulsion with a diameter (mean standard deviation, see Fig.3 and Methods).
2.2.2 Concentrations of lipids and binders
@@
From their size distributions, we can estimate the quantity of surfactant required to cover the surface area of the inner water droplets in the emulsion. The size distribution leads to a mean volume of for the water droplets, while the water volume fraction , with and , yields (see Methods). This allows one to calculate the total oil/water interface in the emulsion and the number of lipids needed to cover it assuming a typical surface coverage per lipid of For the chosen lipids (16:0 Biotinyl PE, ) this leads to a minimum mass concentration of to cover the interface of the inner water droplets. Because the lipids are dispersed in the oil and do not all partition at the oil/water interface, we therefore work at lipid concentrations .
@@
The binders concentration also needs to be adapted to the amount of available lipids on the surface. In order to avoid the presence of unattached binders in the inner water droplets, their concentration must be lower than that of the surface lipids. Moreover, as mentioned above, not all lipids will partition at the oil/water interface, especially in the presence of co-surfactants such as Tween 20. Therefore we work at a ratio of about 1 streptavidin for 2000 lipids. Considering the size of the water droplets in the inverted emulsion, this target value leads to a streptavidin concentration of .
2.2.3 Emulsification through filters
@@
If the material contained in the water droplets is more robust and abundant, the emulsification method can be simply performed by pushing the oil and water mix through a classical PVDF (Polyvinylidene fluoride) filter with a pore size (see Methods). The dead volume with these filters is quite important, which is why it is not adapted to the handling of precious material such as purified proteins, but can be used for classical fluorophores or drugs. For the same formulation (see light pink points in Fig. 3B), we produced an emulsion that is quite comparable to the one obtained with SPG membranes, with a diameter (mean standard deviation). With an alternative formulation containing only lipids, we get a coarser emulsion, with (see blue points in Fig. 3B), which is expected since the surface tension of this formulation should be higher (see next section for values).
@@ 2.3 Tuning of the interfacial tension
In order to use oil droplets as stress sensors in a medium of modulus , their interfacial tension must be low enough and is therefore a critical parameter of our design. Indeed, as mentioned in the introduction, if the size of the oil droplet is much smaller than the elastocapillary length , the capillary stresses will dominate and the oil droplet will behave as a solid inclusion in the tissue. Conversely, if it is larger or on the order of the elastocapillary length, the local elastic stresses will allow to deform the oil droplet that can then serve as a stress sensor (see Fig.4)[37].
@@
Therefore the radius of the oil droplet should satisfy . In soft developing tissues or biopolymers such as extracellular matrices, the material Young’s modulus typically falls in the range [38]. The oil droplet size is constrained by the one of the considered biological tissue and the typical length scale of the process being probed. Typically, oil droplets of radius are usually small enough to probe cellular scale processes in tissues and large enough to be imaged with a sufficient resolution. We therefore use this minimal size to deduce the required maximal interfacial tension for observable oil droplet deformations, which yields:
@@
To assess if such values are accessible in our system, we first tested the effect of lipids alone on interfacial tension. To this end, we measured the interfacial tension between an oil containing the lipids and an aqueous phase through a rising drop method (see Fig. 5 and Methods). Importantly, we also measured the interfacial tension between an inverted emulsion prepared as mentioned above and a continuous aqueous phase, which matches more accurately the configuration of our system (see Fig. 5C-E). Note that the surfactants require some time to adsorb at the interface of the rising drop. Therefore we track the interfacial tension over time in order to determine its final equilibrium value (see Fig. 5E and Methods).
@@
As the concentration of lipids is increased from to in the oil, the interfacial tension decreases. Notably, the same concentrations of lipids lead to similar interfacial tensions with a rising drop made of an inverted emulsion (see Fig. 6), thus validating our system for stress measurements in vivo. However, in both cases, the tension does not go below , even at lipids. Yet, at such lipid concentrations we expect the inner water droplets to be quite stable [31] and to hinder the targeted self-functionalizing features of the sensor.
@@
To overcome this limitation, the use of Tween 20 as a co-surfactant is critical. Indeed, it is a synthetic surfactant that is largely used to stabilize direct oil in water emulsions, with a HLB index of 16.7. Therefore, we expect that the addition of Tween 20 should preferentially lower the oil/water interface at the outer surface of the sensor (positive curvature), while having a more marginal effect on the stability of the inner water droplets whose negative curvature does not favor its insertion. To test the effect of Tween 20 on surface tension, we thus measured the oil/water interfacial tension for oils containing a fixed concentration of lipids at and Tween 20 concentrations up to (see Fig. 6).
@@
Strikingly, an addition of Tween 20 already lowers the interfacial tension down to , for both the oil and inverted emulsion systems. Increasing the Tween 20 concentration up to only marginally lowers the tension down to and could again hinder the destabilization process. However, in both cases, this value of surface tension is in principle low enough to measure stresses inside soft tissues as discussed above. In the following, we explore the efficiency of the inverted emulsion destabilization, and the resulting self-functionalization, as a function of surfactant concentrations and composition.